a collage planning committee consists of 3 freshmen, 4 sophomores, 5 junious and 2 seniors. a sub-committe of 4consists of 1 individual from each class. how many different sub committees are possible?
Answer by ewatrrr(24785)

 
Hi
a collage planning committee consists of 3 freshmen, 4 sophomores, 5 junious and 2 seniors.
sub-committe of 4consists of 1 individual from each class
how many different sub committees are possible?
3C1*4C1*5C1*2C1 = 3*4*5*2 = 120 ways